A body of mass 25 kg moving at 3 m per second on a rough horizontal floor is brought to rest after sliding through a distance of 2.5 m on the floor. Calculate the coefficient of sliding friction (g=10.0ms-2)




1
Expert's answer
2022-04-21T13:01:18-0400

"\frac{mv^2}2=\mu mgl,"

"\mu=\frac{v^2}{2gl}=0.18."